# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.3. The case $\Delta=8$
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.28, $m_{3}(v)=0$ and $m_{4}(v)\leq 1$. If $m_{4}(v)=0$, then $v$ is adjacent to two $8(6^{-})$-vertices by Lemma 2.28(b), and so $\mu^{*}(v)\geq-1+3\times\frac{1}{5}+2\times\frac{1}{5}=0$ after $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and $\frac{1}{5}$ from each of its $8(6^{-})$-neighbours by R3. On the other hand, if $m_{4}(v)=1$, then $v$ is adjacent to three $8(6^{-})$-neighbours by Lemma 2.28(c), and so $\mu^{*}(v)\geq-1+2\times\frac{1}{5}+3\times\frac{1}{5}=0$ after $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and $\frac{1}{5}$ from each of its $8(6^{-})$-neighbours by R3.
@@
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)\geq 2$, then $v$ has two $7^{+}$-neighbours by Lemma 2.30, which are clearly different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices as $m_{3}(v)=1$. It follows that $v$ receives $\frac{1}{6}$ from each of its $7^{+}$-neighbours by R4. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
@@
If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices by Lemma 2.31(a). It follows that $v$ receives totally at least $2\times\frac{1}{6}$ from its $7^{+}$-neighbours by R5. In addition, $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}+2\times\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices by Lemma 2.31(b). In each case, $v$ receives totally at least $\frac{1}{2}$ from its $7^{+}$-neighbours by R5. Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{2}+\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $m_{4}(v)=2$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours by Lemma 2.31(c). In each case, $v$ receives totally at least $\frac{2}{3}=\min\{\frac{1}{4}+3\times\frac{1}{6},2\times\frac{1}{4}+\frac{1}{6},3\times\frac{1}{4}\}$ from its $7^{+}$-neighbours by R5. Thus, $\mu^{*}(v)\geq\frac{2}{3}-2\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(2.3). Let $m_{3}(v)=3$. Suppose first that $m_{4}(v)=0$. It then follows from Lemma 2.32(a) that $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or two $8(7^{-})$- and two 8(8)888(8)8 ( 8 )-neighbours or three $8(7^{-})$-neighbours. In each case, $v$ receives totally at least $\frac{5}{6}=\min\{\frac{1}{3}+3\times\frac{1}{6},2\times\frac{1}{3}+\frac{1}{6},2\times\frac{1}{3}+2\times\frac{1}{9},3\times\frac{1}{3}\}$ from its $7^{+}$-neighbours by R6. In addition, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{5}{6}+\frac{1}{5}-3\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Now, we suppose that $m_{4}(v)=1$. By Lemma 2.32(b), $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours. In both cases, $v$ receives totally at least $1=\min\{2\times\frac{1}{3}+2\times\frac{1}{6},3\times\frac{1}{3}\}$ from its $7^{+}$-neighbours by R6. Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(2.4). Let $m_{3}(v)=4$. By Lemma 2.33, $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour or four $8(7^{-})$-neighbours. By applying R7, $v$ receives $\frac{1}{3}$ from each of its $7^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 4\times\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.1). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.35(a), $v$ has two $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. It follows that $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex by R8, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Next we suppose that $m_{4}(v)=1$. By Lemma 2.35(b), $v$ has three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. So, $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex by R8. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.2.1). Let $n_{4}(v)=0$. If $n_{5}(v)=0$, then, by Lemma 2.36(b), $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex, two of which are $7^{+}$-vertices different from 7(7)777(7)7 ( 7 )-vertex. It follows from applying R9 that $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $n_{5}(v)=1$, then, by Lemma 2.36(c), $v$ has either three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. So, $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours by R9. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex by Lemma 2.36(d). By R9(b), $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. Thus, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.2.2). Let $n_{4}(v)=1$. By Lemma 2.36(a), $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. It follows that $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex by R9(b). So, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(4). Let $k=6$. The initial charge of $v$ is $\mu(v)=d(v)-4=2$. If $m_{3}(v)\leq 4$, then $\mu^{*}(v)\geq 2-4\times\frac{1}{3}-6\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a). On the other hand, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $m_{3}(v)=5$. By Lemma 2.37, $v$ has at most four $5(4^{+})$-neighbours. If $v$ has at most three $5(4^{+})$-neighbours, then $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a). Next we assume that $v$ has exactly four $5(4^{+})$-neighbours. If $m_{4}(v)=0$, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2, and so $\mu^{*}(v)\geq 2+\frac{1}{5}-5\times\frac{1}{3}-4\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a). If $m_{4}(v)=1$, then $v$ has two $8(6^{-})$-neighbours by Lemma 2.37, and so $v$ receives $\frac{1}{18}$ from each of its $8(6^{-})$-neighbours by R10. Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{18}-5\times\frac{1}{3}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a).
@@
(5). Let $k=7$. The initial charge of $v$ is $\mu(v)=d(v)-4=3$. Notice first that if $m_{3}(v)\leq 3$, then we have $\mu^{*}(v)\geq 3-3\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its 4(4)444(4)4 ( 4 )-neighbours by R7, and at most $\frac{2}{9}$ to each of its other neighbours by R4-R6 and R8-R9, where we note that $v$ can have at most two 4(4)444(4)4 ( 4 )-neighbours due to $m_{3}(v)\leq 3$. Therefore, we may assume that $4\leq m_{3}(v)\leq 7$.
@@
(5.1). Let $m_{3}(v)=4$. Observe that $v$ can have at most two neighbours consisting of 4(4)444(4)4 ( 4 )- or 5(5)555(5)5 ( 5 )-vertices by Lemmas 2.29(c) and 2.34(b). If $v$ has at most one 4(4)444(4)4 ( 4 )-neighbour, then $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-\frac{1}{3}-6\times\frac{2}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its 4(4)444(4)4 ( 4 )-neighbour by R7, and at most $\frac{2}{9}$ to each of its other neighbours by R4-R6 and R8-R9. If $v$ has exactly two 4(4)444(4)4 ( 4 )-neighbours, then $v$ is not adjacent to any 5(5)555(5)5 ( 5 )-vertex as stated earlier, and so $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its 4(4)444(4)4 ( 4 )-neighbours by R7, and at most $\frac{1}{6}$ to each of its other neighbours by R4-R6 and R8.
@@
(5.2). Let $m_{3}(v)=5$. By Lemmas 2.29(c) and 2.34(b), $v$ has at most two 4(4)444(4)4 ( 4 )-neighbours. If $v$ has no 4(4)444(4)4 ( 4 )-neighbour, then $v$ is adjacent to at most three 5(5)555(5)5 ( 5 )-neighbour by Lemma 2.34(b), and so $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{2}{9}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), and at most $\frac{1}{6}$ to each of its other neighbours by R4-R6 and R8. If $v$ has exactly one 4(4)444(4)4 ( 4 )-neighbour, then $v$ is adjacent to at most two 5(5)555(5)5 ( 5 )-neighbours by Lemmas 2.29(c) and 2.34(b). Moreover, $v$ has at least two neighbours $v_{i}$ and $v_{j}$ that are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.29(c), since $v$ has a 4(4)444(4)4 ( 4 )-neighbour. Clearly, neither $v_{i}$ nor $v_{j}$ receives any charge from $v$. Therefore, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-\frac{1}{3}-2\times\frac{2}{9}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its 4(4)444(4)4 ( 4 )-neighbour by R7, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from 4(4)444(4)4 ( 4 )-vertex by R4-R6 and R8. If $v$ has exactly two 4(4)444(4)4 ( 4 )-neighbours, then $v$ is not adjacent to any 5(5)555(5)5 ( 5 )-vertices by Lemma 2.29(c). Moreover, $v$ has at least three neighbors $v_{i}$, $v_{j}$, and $v_{k}$ that are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.29(c), since $v$ has two 4(4)444(4)4 ( 4 )-neighbors. Clearly, none of $v_{i}$, $v_{j}$, or $v_{k}$ receives any charge from $v$. Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{3}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its 4(4)444(4)4 ( 4 )-neighbours by R7, and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from 4(4)444(4)4 ( 4 )-vertex by R4-R6 and R8.

### v3

@@ 2.3. The case $\Delta=8$
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.30, $m_{3}(v)=0$ and $m_{4}(v)\leq 1$. If $m_{4}(v)=0$, then $v$ is adjacent to two $8(6^{-})$-vertices by Lemma 2.30(b), and if $m_{4}(v)=1$, then $v$ is adjacent to three $8(6^{-})$-neighbours by Lemma 2.30(c). Thus $\mu^{*}(v)\geq-1+3\times\frac{1}{5}+2\times\frac{1}{5}=0$ after $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and $\frac{1}{5}$ from each of its $8(6^{-})$-neighbours by R3.
@@
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)\geq 2$, then $v$ has two $7^{+}$-neighbours by Lemma 2.32, which are clearly different from $7(7)$- and $8(8)$-vertices as $m_{3}(v)=1$. It follows that $v$ receives $\frac{1}{6}$ from each of its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(a). Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
@@
If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours different from $7(7)$- and $8(8)$-vertices by Lemma 2.33(a). It follows that $v$ receives totally at least $2\times\frac{1}{6}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(b). In addition, $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}+2\times\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours different from $7(7)$- and $8(8)$-vertices by Lemma 2.33(b). In each case, $v$ receives totally at least $\frac{1}{2}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(b). Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{2}+\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $m_{4}(v)=2$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours by Lemma 2.33(c). In each case, $v$ receives totally at least $\frac{2}{3}=\min\{\frac{1}{4}+3\times\frac{1}{6},2\times\frac{1}{4}+\frac{1}{6},3\times\frac{1}{4}\}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(b). Thus, $\mu^{*}(v)\geq\frac{2}{3}-2\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(2.3). Let $m_{3}(v)=3$. Suppose first that $m_{4}(v)=0$. It then follows from Lemma 2.34(a) that $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or two $8(7^{-})$- and two $8(8)$-neighbours or three $8(7^{-})$-neighbours. In each case, $v$ receives totally at least $\frac{5}{6}=\min\{\frac{1}{3}+3\times\frac{1}{6},2\times\frac{1}{3}+\frac{1}{6},2\times\frac{1}{3}+2\times\frac{1}{9},3\times\frac{1}{3}\}$ from its $7(6^{-})$- and $8$-neighbours by R7(a), R9(c), R10(a). In addition, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{5}{6}+\frac{1}{5}-3\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Now, we suppose that $m_{4}(v)=1$. By Lemma 2.34(b), $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours. In both cases, $v$ receives totally at least $1=\min\{2\times\frac{1}{3}+2\times\frac{1}{6},3\times\frac{1}{3}\}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(c). Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(2.4). Let $m_{3}(v)=4$. By Lemma 2.35, $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour or four $8(7^{-})$-neighbours. By applying R4 and R9(c), $v$ receives $\frac{1}{3}$ from each of its $7(5^{-})$- and $8(7^{-})$-neighbours. Thus, $\mu^{*}(v)\geq 4\times\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.1). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.36(a), $v$ has two $6^{+}$-neighbours different from $6(6)$-vertex. It follows that $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from $6(6)$-vertex by R6-R10, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Next we suppose that $m_{4}(v)=1$. By Lemma 2.36(b), $v$ has three $6^{+}$-neighbours different from $6(6)$-vertex. So, $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from $6(6)$-vertex by R6-R10. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.2.1). Let $n_{4}(v)=0$. If $n_{5}(v)=0$, then, by Lemma 2.38(b), $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from $6(6)$-vertex, two of which are $7^{+}$-vertices different from $7(7)$-vertex. It follows from applying R6-R10 that $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $n_{5}(v)=1$, then, by Lemma 2.38(c), $v$ has either three $7^{+}$-neighbours different from $7(7)$-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from $7(7)$-vertex. So, $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours by R6-R10. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
If $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from $7(7)$-vertex by Lemma 2.38(d). It follows that $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from $7(7)$-vertex by R7(c), R9(e), R10(b). Thus, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.2.2). Let $n_{4}(v)=1$. By Lemma 2.38(a), $v$ has three $7^{+}$-neighbours different from $7(7)$-vertex. It follows that $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from $7(7)$-vertex by R7(c), R9(e), R10(b). So, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(4). Let $k=6$. The initial charge of $v$ is $\mu(v)=d(v)-4=2$. If $m_{3}(v)\leq 4$, then $\mu^{*}(v)\geq 2-4\times\frac{1}{3}-6\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6. On the other hand, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $m_{3}(v)=5$. By Lemma 2.39, $v$ has at most four $5(4^{+})$-neighbours. If $v$ has at most three $5(4^{+})$-neighbours, then $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6. Next we assume that $v$ has exactly four $5(4^{+})$-neighbours. If $m_{4}(v)=0$, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2, and so $\mu^{*}(v)\geq 2+\frac{1}{5}-5\times\frac{1}{3}-4\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6. If $m_{4}(v)=1$, then $v$ has two $8(6^{-})$-neighbours by Lemma 2.39, and so $v$ receives $\frac{1}{18}$ from each of its $8(6^{-})$-neighbours by R5. Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{18}-5\times\frac{1}{3}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6.
@@
(5). Let $k=7$. The initial charge of $v$ is $\mu(v)=d(v)-4=3$. Notice first that if $m_{3}(v)\leq 3$, then we have $\mu^{*}(v)\geq 3-3\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(4)$-neighbours by R4, and at most $\frac{2}{9}$ to each of its other neighbours by R7, where we note that $v$ can have at most two $4(4)$-neighbours due to $m_{3}(v)\leq 3$. Therefore, we may assume that $4\leq m_{3}(v)\leq 7$.
@@
(5.1). Let $m_{3}(v)=4$. Observe that $v$ can have at most two neighbours consisting of $4(4)$- or $5(5)$-vertices by Lemmas 2.31(c) and 2.37(b). If $v$ has at most one $4(4)$-neighbour, then $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-\frac{1}{3}-6\times\frac{2}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its $4(4)$-neighbour by R4, and at most $\frac{2}{9}$ to each of its other neighbours by R7. If $v$ has exactly two $4(4)$-neighbours, then $v$ is not adjacent to any $5(5)$-vertex as stated earlier, and so $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(4)$-neighbours by R4, and at most $\frac{1}{6}$ to each of its other neighbours by R7(a)-(b).
@@
(5.2). Let $m_{3}(v)=5$. By Lemma 2.31(c), $v$ has at most two $4(4)$-neighbours. If $v$ has no $4(4)$-neighbour, then $v$ is adjacent to at most three $5(5)$-neighbour by Lemma 2.37(b), and so $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{2}{9}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{6}$ to each of its other neighbours by R7(a)-(b). If $v$ has exactly one $4(4)$-neighbour, say $v_{i}$, then $v$ is adjacent to at most two $5(5)$-neighbours by Lemmas 2.31(c) and 2.37(b). Moreover, the neighbours of $v_{i}$ belonging to $G[N(v)]$, say $v_{j}$ and $v_{k}$, are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.31(c). Clearly, neither $v_{j}$ nor $v_{k}$ receives any charge from $v$. Therefore, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-\frac{1}{3}-2\times\frac{2}{9}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its $4(4)$-neighbour by R4, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from $4(4)$-vertex by R7(a)-(b). If $v$ has exactly two $4(4)$-neighbours, then $v$ is not adjacent to any $5(5)$-vertices by Lemma 2.31(c). Moreover, similarly as above, $v$ has at least three neighbours $v_{i}$, $v_{j}$, and $v_{k}$ that are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.31(c), since $v$ has two $4(4)$-neighbors. Clearly, none of $v_{i}$, $v_{j}$, or $v_{k}$ receives any charge from $v$. Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{3}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(4)$-neighbours by R4, and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from $4(4)$-vertex by R7(a)-(b).